OpenStudy (anonymous):

Help pleaseeeeeeeeeee =*( Which one is greater or are they equal? A. The product of BE and AC B. The product of BC and AD

OpenStudy (anonymous):



OpenStudy (anonymous):

where is point D?

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

BC and AD are equal

OpenStudy (anonymous):

how

OpenStudy (anonymous):

i mean BE and AC are equal

OpenStudy (anonymous):

how do u know that

OpenStudy (anonymous):

its fine and if these are two separate questions then part A of it that line BE is equal to line AC, part B of it is that line BC is greater than line AD because of line AD is going at a different angle to BC and thus making line AD shorter than line BC but if it were line to BE DC and line AD then they would be reversed and line AD will be slightly longer and line BE is going done a equilateral triangle meaning if you attach another triangle with the same lengths under it line AC will be equal

OpenStudy (anonymous):

yea that y i said never mind on BC and AD

OpenStudy (anonymous):

slightly confused so a and b are equal?

OpenStudy (anonymous):

no, line AC and line BE are equal but line BC is greater than line AD

OpenStudy (anonymous):

the answer says A and B are equal...

OpenStudy (anonymous):

ohh, i get it know, it referring to where the line intercept and what angle they make, so in this case A is equal to B because the angles of both are 90 degrees

OpenStudy (anonymous):

Man i'm so confused....... can you start from the beginning and break it down so i can understand better

OpenStudy (anonymous):

read the lines, now look at where the line intercept

OpenStudy (anonymous):

the question is asking if the angle the intercepted lines make are equal or uneven

OpenStudy (anonymous):

no it's asking which one is greater or are they equal. But go on

OpenStudy (anonymous):

they both form right angles as shown in the picture

OpenStudy (anonymous):

so the angle from A and the angle from B both are right angles making them equal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Use the formula for area of triangle In \[\triangle ABC\] We have consider AC as base and BE as height, we have \[A_{1}=\frac{1}{2}\times AC \times BE\] Now consider BC as base and AD as height we have \[A_{2}=\frac{1}{2} \times BC \times AD\] We are measuring area of the same triangle, so both areas should be equal \[A_{1}=A_{2} \implies \frac{1}{2} \times AC \times BE=\frac{1}{2}\times BC \times AD\] \[\implies AC \times BE=BC \times AD\]

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

because it's the same triangle they're equal no matter what? But they're asking for the product... will the measurements still be equal?

random231 (random231):

@Nishant_Garg is right!

OpenStudy (anonymous):

No, what I'm saying is I've measured area of triangle ABC in 2 different ways, since I'm only measuring area of ABC, regardless of how I measure, as long as my method is correct, both area should be equal as we are only talking about the same triangle ABC

OpenStudy (anonymous):

gotcha! thanks a million!!!